# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 4.3.1 Type I vertex operator
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})=\mathrm{Sym}_{\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}}\left(\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}\frac{\theta(x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(t_{1}x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\prod_{b\in\lambda^{\prime\prime}}\left(-\frac{\theta(\hbar u_{1}/x_{b}^{\prime\prime})}{\theta(x^{\prime\prime}_{b}/u_{1})}\right)\right.{}$
×$\hskip 170.71652pt\qquad\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z}\hbar^{-1})\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z})\Biggr{)}$
Comparing this with (3.10), one finds that $\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ coincides with the hatted version of the elliptic stable envelope for $\mathrm{E}_{T}({\cal M}(n^{\prime}+n^{\prime\prime},2))$ given in (3.5). We thus regard the composition $\Phi_{\lambda^{\prime}}(u_{1})\Phi_{\lambda^{\prime\prime}}(u_{2})$ as a vertex operator for ${\cal M}(n^{\prime}+n^{\prime\prime},2)$.
By repeating such consideration, one obtains a vertex operator for $\bigoplus_{n}\mathrm{E}_{T}({\cal M}(n,r))$ by the following composition.
$({\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id}\widetilde{\otimes}\Phi(u_{1}))\circ\cdots({\rm id}\widetilde{\otimes}\Phi(u_{r-1}))\circ\Phi(u_{r})\ :\ {\cal F}^{(1,N+r)}_{v}\ \to\ {\cal F}^{(0,1)}_{u_{r}}\widetilde{\otimes}\cdots\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{1}}\widetilde{\otimes}{\cal F}^{(1,N)}_{(-)^{r}v/u_{1}\cdots u_{r}}.$
@@
$({\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id}\widetilde{\otimes}\Phi(u_{1}))\circ\cdots({\rm id}\widetilde{\otimes}\Phi(u_{r-1}))\circ\Phi(u_{r}){}$
$=\sum_{n\in{\mathbb{Z}}_{\geq 0}}\sum_{{\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})\atop|{\boldsymbol{\lambda}}|=n}{|\lambda^{(r)}\rangle}_{u_{r}}\widetilde{\otimes}\cdots\widetilde{\otimes}{|\lambda^{(1)}\rangle}_{u_{1}}\widetilde{\otimes}\ \Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r}).$ (4.23)
@@
$\Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})=\Phi_{\lambda^{(1)}}(u_{1})\cdots\Phi_{\lambda^{(r)}}(u_{r}){}$
$=\int_{{\cal C}^{r}}\prod_{a\in{\boldsymbol{\lambda}}}dx_{a}:\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):\ \Phi_{\emptyset}(u_{1})\cdots\Phi_{\emptyset}(u_{r})\ \prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z}),{}$
where $\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})$ is the hatted version of the elliptic stable envelope for $\mathrm{E}_{T}({\cal M}(n,r))$ given in (3.8) satisfying the shuffle product formula (3.10). We hence regard $\Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})$ as the type I vertex operator for ${\cal M}(n,r)$. Then the total one (4.23) is the vertex operator for $\bigoplus_{n}{\cal M}(n,r)$.
4.3.2 Type II dual vertex operator
@@
$\Psi^{*}(u_{r})\circ(\Psi^{*}(u_{r-1})\widetilde{\otimes}{\rm id})\circ\cdots\circ(\Psi^{*}(u_{1})\widetilde{\otimes}{\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id})\ :\ {\cal F}^{(1,N)}_{v}\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{1}}\widetilde{\otimes}\cdots\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{r}}\ \to\ {\cal F}^{(1,N+r)}_{(-)^{r}vu_{1}\cdots u_{r}}.{}$
@@
$\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})\xi$ $=$ $\Psi^{*}(u_{r})\circ(\Psi^{*}(u_{r-1})\widetilde{\otimes}{\rm id})\circ\cdots\circ(\Psi^{*}(u_{1})\widetilde{\otimes}{\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id})(\xi\widetilde{\otimes}{|\lambda^{(1)}\rangle}_{u_{1}}\widetilde{\otimes}\cdots\widetilde{\otimes}{|\lambda^{(r)}\rangle}_{u_{r}})$
$=$ $\Psi^{*}_{\lambda^{(r)}}(u_{r})\Psi^{*}_{\lambda^{(r-1)}}(u_{r-1})\cdots\Psi^{*}_{\lambda^{(1)}}(u_{1})\xi,\quad\qquad\forall\xi\in{\cal F}^{(1,N)}_{v}.$
@@
$\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r}){}$
$=\int_{{\cal C}^{*r}}\prod_{a\in{\boldsymbol{\lambda}}}dx_{a}\ \widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})\ :\prod_{a\in{\boldsymbol{\lambda}}}x^{+}(x_{a}):\ \Psi^{*}_{\emptyset}(u_{r})\cdots\Psi^{*}_{\emptyset}(u_{1}){}$
×$\qquad\qquad\qquad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym}.$ (4.26)
Here $\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ is given in (3.12). We hence regard $\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})$ as the type II dual vertex operator for ${\cal M}(n,r)$. The total one (4.25) is the vertex operator for $\bigoplus_{n}{\cal M}(n,r)$.
The shuffle product formula for $\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ is given by (3.13). This is reproduced by considering a composition $\Psi^{*}_{{\boldsymbol{\lambda}}^{\prime\prime}}(u^{\prime\prime}_{1},\cdots,u^{\prime\prime}_{r^{\prime\prime}})\Psi^{*}_{{\boldsymbol{\lambda}}^{\prime}}(u^{\prime}_{1},\cdots,u^{\prime}_{r^{\prime}})$ and making the following arrangements of the operators in the integrand.
@@
    Move $\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C}^{\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}^{*-1})$ to the left of all operators.
@@
    Move $:\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):$ to the left of $\Psi^{*}_{\emptyset}(u^{\prime\prime}_{r^{\prime\prime}})\cdots\Psi^{*}_{\emptyset}(u^{\prime\prime}_{1})$ by using the formula (4.15).
@@
    Make $:\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}x^{+}(x^{\prime\prime}_{b})::\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):$ totally normal ordered product by the formula
    $:\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}x^{+}(x^{\prime\prime}_{b})::\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):=\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}<x^{+}(x^{\prime\prime}_{b})x^{+}(x^{\prime}_{a})>\ :\prod_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}x^{+}(x_{a}):.$
@@
    Divide $<x^{+}(x^{\prime\prime}_{b})x^{+}(x^{\prime}_{a})>$ into the symmetric part (4.3) and the non-symmetric part (4.5).
@@
5 The $\mathrm{K}$-Theoretic Vertex functions for ${\cal M}(n,r)$
In this section, we calculate vacuum expectation values of the vertex operators constructed in the last section and show that they give $\mathrm{K}$-theoretic vertex functions for ${\cal M}(n,r)$.
5.1 The type I case
@@
$\Phi_{\lambda}(u)=\prod_{a\in\lambda}\left(\int_{0}^{\varphi^{\mu}_{\varsigma(a)}}d_{p}x_{a}\right):\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u)\prod_{a,b\in\lambda\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ {}$
×$\hskip 170.71652pt\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}).$ (5.1)
Let ${|0\rangle}_{v}^{(1,N)}$ be the vacuum state (2.25) in ${\cal F}^{(1,N)}_{v}$ and ${}^{(1,N)}{}_{v}\langle 0|$ be the dual state satisfying
${}^{(1,N)}{}_{v}\langle 0|{|0\rangle}_{v}^{(1,N)}=1.$
@@
$:\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u)=\prod_{a\in\lambda}\left(-\frac{\hbar^{1/2}u}{x_{a}}\right)\frac{(pu/x_{a};p)_{\infty}}{(\hbar u/x_{a};p)_{\infty}}:\prod_{a\in\lambda}x^{-}(x_{a})\cdot\Phi_{\emptyset}(u):.$
@@
$\prod_{a\in\lambda}e^{-h}\left(\hbar^{1/2}/x_{a}\right)^{-c^{\perp}}\cdot(-u)^{-\alpha}e^{-\Lambda_{0}^{\perp}}=\prod_{a\in\lambda}\left(-\frac{\hbar^{1/2}u}{x_{a}}\right)\cdot(-u)^{-\alpha}e^{-\Lambda_{0}^{\perp}}\prod_{a\in\lambda}e^{-h}\left(\hbar^{1/2}/x_{a}\right)^{-c^{\perp}}.$
@@
${}^{(1,N-1)}{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}{}$
@@
×$\qquad\qquad\qquad\qquad\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(pt_{1}^{-1}x_{a}/x_{b};p)_{\infty}(\hbar x_{a}/x_{b};p)_{\infty}}{(px_{a}/x_{b};p)_{\infty}(t_{2}x_{a}/x_{b};p)_{\infty}}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}).$ (5.2)
@@
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\mu}_{\varsigma(a)}p^{d_{a}}\atop(a\in\lambda)}=\mathfrak{z}^{-\sum_{a\in\lambda}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\mu}_{\varsigma(a)}\atop(a\in\lambda)}=\mathfrak{z}^{-\sum_{a\in\lambda}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{\mu}.$

### v2

@@ 4.3.1. The type I vertex operator
$\hskip-28.45274pt\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$
$\hskip-28.45274pt=\mathrm{Sym}_{\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}}\left(\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}\frac{\theta(x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(t_{1}x^{\prime\prime}_{b}/x^{\prime}_{a})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\prod_{b\in\lambda^{\prime\prime}}\left(-\frac{\theta(\hbar u_{1}/x_{b}^{\prime\prime})}{\theta(x^{\prime\prime}_{b}/u_{1})}\right)\right$.{}
$\hskip 28.45274pt\qquad\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z}\hbar^{-1})\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z})\Biggr{)}.$
Comparing this with (3.9), one finds that $\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ coincides with the hatted version of the elliptic stable envelope for $\mathrm{E}_{T}({\mathcal{M}}(n^{\prime}+n^{\prime\prime},2))$ given in (3.4). We thus regard the composition $\Phi_{\lambda^{\prime}}(u_{1})\Phi_{\lambda^{\prime\prime}}(u_{2})$ as a vertex operator for ${\mathcal{M}}(n^{\prime}+n^{\prime\prime},2)$.
By repeating such consideration, one obtains a vertex operator for $\bigoplus_{n}{\mathcal{M}}(n,r)$ by the following composition.
$\hskip-28.45274pt({\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id}\widetilde{\otimes}\Phi(u_{1}))\circ\cdots({\rm id}\widetilde{\otimes}\Phi(u_{r-1}))\circ\Phi(u_{r})\$
$\qquad:\ {\mathcal{F}}^{(1,N)}_{v}\ \to\ {\mathcal{F}}^{(0,-1)}_{u_{r}}\widetilde{\otimes}\cdots\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{1}}\widetilde{\otimes}{\mathcal{F}}^{(1,N+r)}_{(-)^{r}v/u_{1}\cdots u_{r}}$.
@@
$\hskip 5.69054pt({\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id}\widetilde{\otimes}\Phi(u_{1}))\circ\cdots({\rm id}\widetilde{\otimes}\Phi(u_{r-1}))\circ\Phi(u_{r})$ (4.19)
$=\hskip-5.69054pt\sum_{n\in{\mathbb{Z}}_{\geq 0}}\sum_{{\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})\atop|{\boldsymbol{\lambda}}|=n}\hskip-5.69054pt{|\lambda^{(r)}\rangle}_{u_{r}}\widetilde{\otimes}\cdots\widetilde{\otimes}{|\lambda^{(1)}\rangle}_{u_{1}}\widetilde{\otimes}\ \Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r}).{}$
@@
$\Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})=\Phi_{\lambda^{(1)}}(u_{1})\cdots\Phi_{\lambda^{(r)}}(u_{r})$ (4.20)
$=\int_{{\mathcal{C}}^{r}}\prod_{a\in{\boldsymbol{\lambda}}}dx_{a}:\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):\ \Phi_{\emptyset}(u_{1})\cdots\Phi_{\emptyset}(u_{r}){}$
$\qquad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z}),{}$
where $\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})$ is the hatted version of the elliptic stable envelope for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$ given in (3.7) satisfying the shuffle product formula (3.9). We hence regard $\Phi_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})$ as the type I vertex operator for ${\mathcal{M}}(n,r)$. Then the total one (4.19) is the vertex operator for $\bigoplus_{n}{\mathcal{M}}(n,r)$.
4.3.2. The type II dual vertex operator
@@
$\Psi^{*}(u_{r})\circ(\Psi^{*}(u_{r-1})\widetilde{\otimes}{\rm id})\circ\cdots\circ(\Psi^{*}(u_{1})\widetilde{\otimes}{\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id})$ (4.21)
$\qquad:\ {\mathcal{F}}^{(1,N)}_{v}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{1}}\widetilde{\otimes}\cdots\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{r}}\ \to\ {\mathcal{F}}^{(1,N-r)}_{(-)^{r}vu_{1}\cdots u_{r}}$.{}
@@
$\hskip-28.45274pt\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})\xi{}$
$\hskip-28.45274pt=\Psi^{*}(u_{r})\circ(\Psi^{*}(u_{r-1})\widetilde{\otimes}{\rm id})\circ\cdots\circ(\Psi^{*}(u_{1})\widetilde{\otimes}{\rm id}\widetilde{\otimes}\cdots\widetilde{\otimes}{\rm id}){}$
$\qquad\qquad\qquad\cdot(\xi\widetilde{\otimes}{|\lambda^{(1)}\rangle}_{u_{1}}\widetilde{\otimes}\cdots\widetilde{\otimes}{|\lambda^{(r)}\rangle}_{u_{r}})$
$\hskip-28.45274pt=\Psi^{*}_{\lambda^{(r)}}(u_{r})\Psi^{*}_{\lambda^{(r-1)}}(u_{r-1})\cdots\Psi^{*}_{\lambda^{(1)}}(u_{1})\xi,\quad\forall\xi\in{\mathcal{F}}^{(1,N)}_{v}$.
@@
$\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})$ (4.22)
$=\int_{{\mathcal{C}}^{*r}}\prod_{a\in{\boldsymbol{\lambda}}}dx_{a}\ \widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})\ :\prod_{a\in{\boldsymbol{\lambda}}}x^{+}(x_{a}):\ {}$
$\qquad\times\Psi^{*}_{\emptyset}(u_{r})\cdots\Psi^{*}_{\emptyset}(u_{1})\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}>\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym}.{}$
Here $\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ is given in (3.11) with $\overline{\mathfrak{C}}:|u_{1}|<\cdots<|u_{r}|$. We hence regard $\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r})$ as the type II dual vertex operator for ${\mathcal{M}}(n,r)$. The total one (4.21) is the vertex operator for $\bigoplus_{n}{\mathcal{M}}(n,r)$.
The shuffle product formula for $\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}({\boldsymbol{\lambda}};\mathfrak{z}^{*-1})$ is given by (3.12). This is reproduced by considering a composition
$\Psi^{*}_{{\boldsymbol{\lambda}}^{\prime\prime}}(u^{\prime\prime}_{1},\cdots,u^{\prime\prime}_{r^{\prime\prime}})\Psi^{*}_{{\boldsymbol{\lambda}}^{\prime}}(u^{\prime}_{1},\cdots,u^{\prime}_{r^{\prime}})$ and making the following arrangements of the operators in the integrand.
@@
    Move $\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}}^{\prime},T^{1/2}_{opp}}({\boldsymbol{\lambda}}^{\prime};\mathfrak{z}^{*-1})$ to the left of all operators.
@@
    Move $\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):$ to the left of $\Psi^{*}_{\emptyset}(u^{\prime\prime}_{r^{\prime\prime}})\cdots\Psi^{*}_{\emptyset}(u^{\prime\prime}_{1})$ by using the formula (4.11).
@@
    Make $\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}x^{+}(x^{\prime\prime}_{b})::\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):$ totally normal ordered product by the formula
    $:\prod_{b\in{\boldsymbol{\lambda}}^{\prime\prime}}x^{+}(x^{\prime\prime}_{b})::\prod_{a\in{\boldsymbol{\lambda}}^{\prime}}x^{+}(x^{\prime}_{a}):$
    $=\prod_{a\in{\boldsymbol{\lambda}}^{\prime},b\in{\boldsymbol{\lambda}}^{\prime\prime}}<x^{+}(x^{\prime\prime}_{b})x^{+}(x^{\prime}_{a})>\ :\prod_{a\in({\boldsymbol{\lambda}}^{\prime},{\boldsymbol{\lambda}}^{\prime\prime})}x^{+}(x_{a}):.$
@@
    Divide $x^{+}(x^{\prime\prime}_{b})x^{+}(x^{\prime}_{a})>$ into the symmetric part (4.3) and the non-symmetric part (4.5).
@@
5. The $\mathrm{K}$-Theoretic Vertex Functions for ${\mathcal{M}}(n,r)$
In this section, we calculate vacuum expectation values of the vertex operators constructed in the last section and show that they give $\mathrm{K}$-theoretic vertex functions for ${\mathcal{M}}(n,r)$.
5.1. The type I case
@@
$\Phi_{\lambda}(u)=\prod_{a\in\lambda}\left(\int_{0}^{\varphi^{\mu}_{\varsigma(a)}}d_{p}x_{a}\right):\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u)$ (5.1)
$\hskip 28.45274pt\times\prod_{a,b\in\lambda\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}).{}$
[…]
${}^{(1,N)}\hskip-4.83691pt{}_{v}\langle 0|{|0\rangle}_{v}^{(1,N)}=1.$
@@
$\hskip-28.45274pt:\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u)=\prod_{a\in\lambda}\left(-\frac{\hbar^{1/2}u}{x_{a}}\right)\frac{(pu/x_{a};p)_{\infty}}{(\hbar u/x_{a};p)_{\infty}}:\prod_{a\in\lambda}x^{-}(x_{a})\cdot\Phi_{\emptyset}(u):.$
@@
$\prod_{a\in\lambda}e^{-h}\left(\hbar^{1/2}/x_{a}\right)^{c}\cdot(-u)^{-\alpha}e^{\Lambda_{0}}$
$=\prod_{a\in\lambda}\left(-\frac{\hbar^{1/2}u}{x_{a}}\right)\times(-u)^{-\alpha}e^{\Lambda_{0}}\prod_{a\in\lambda}e^{-h}\left(\hbar^{1/2}/x_{a}\right)^{c}.$
@@
${}^{(1,N-1)}\hskip-19.91684pt{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}$ (5.2)
@@
$\quad\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(pt_{1}^{-1}x_{a}/x_{b};p)_{\infty}(\hbar x_{a}/x_{b};p)_{\infty}}{(px_{a}/x_{b};p)_{\infty}(t_{2}x_{a}/x_{b};p)_{\infty}}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}).{}$
@@
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\mu}_{\varsigma(a)}p^{d_{a}}\atop(a\in\lambda)}$ $=$ $\mathfrak{z}^{-\sum_{a\in\lambda}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\mu}_{\varsigma(a)}\atop(a\in\lambda)}$
$=$ $\mathfrak{z}^{-\sum_{a\in\lambda}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{\mu}.$